1. If the distribution of income in country Cis (1, 2, 2, 3, 5), and in country D its (1, 1, 2, 3, 5). The poverty line
Answer #1

ans a) AIS = TPG/ H

where TPG is total poverty gap

H is head count( no of people below poverty line

Yp is poverty line income

Y is obsevations of income

for C,

TPG= Y<Yp = (2.5-1)+(2.5-2) +( 2.5-2)= 2.5

H= 3

AIS=TPG/H = 2.5/3 =0.83

for D

TPG= Y<Yp = (2.5-1)+(2.5-1) +( 2.5-2) = 3.5

H= 3

AIS = 3.5/3 = 1.17

according to AIS, country "D" is more poor

b) head count index= H/N

where H= head count

N= total population = population size

for C,

head count index= 3/5

for D,

head count index= 3/5

according to head count index , both are equally poor

c)

APG= TPG/ N

for C,

APG=2.5 /5= 0.5

for D

APG =3.5/5 =0.7

according to APG, D is more poor

d) FGT= 1/N \sum_{i=1}^{H}(\frac{Yp-Y}{Yp})^{a}

here a= 2( given) here all Y with Y<Yp will be used

for C,

FGT = 1/ 5 ( (( 2.5 -1)/2.5)2+(( 2.5 -2)/2.5)2+(( 2.5 -2)/2.5)2​​​​​​​) =1/5 (0.36 + 0.04 +0.04) = 0.44/5= 0.088

for D,

FGT =1/ 5 ( (( 2.5 -1)/2.5)2+(( 2.5 -1)/2.5)2​​​​​​​+(( 2.5 -2)/2.5)2​​​​​​​) = 1/5 ( 0.36+ 0.36+ 0.04) =0.76/5 = 0.155